OpenStudy (anonymous):

Factoring Quadratic Expressions: x2 + 9x + 18

OpenStudy (anonymous):

(x+3)(x+6)=0

OpenStudy (anonymous):

thanks! now i need to know how you got that, i have the answers, i just don't understand how to solve the problem.

OpenStudy (anonymous):

Markovian: it would just be (x+3)(x+6)

OpenStudy (anonymous):

We can see that x^2 has no coefficient, therefore we can have x's by themselves on the outside. 6+3=9, which is the x term and 6*3=18 which is the non-x term. You can always FOIL it back to check the answer to get x*x+6x+3x+18=x^2+9x+18.

OpenStudy (anonymous):

And yes it would just be (x+3)(x+6).

OpenStudy (anonymous):

Well I just gave you a medal, thanks so much for explaining it. i get it!